Suppose that a person throws a ball upward at 15 m/s while standing on the edge of a cliff so that the ball can fall to the base of the cliff 50 m below.



a) How long does it take the ball to reach the base of the cliff?



b) What is the total distance covered by the ball.




1
Expert's answer
2022-02-07T08:42:40-0500

a) We can calculate the time with the equation


"h= h_0+V_0t-\frac{gt^2}{2}= 0
\\ t= \frac{-V_0 ±\sqrt{V_0^2+2gh_0 }}{-g} 
\\ t= \frac{V_0}{g} \mp \sqrt{ (\frac{V_0}{g})^2+\frac{2h_0}{g} }
\\ t= (\frac{V_0}{g})\Big(1\mp \sqrt{ 1+\frac{2gh_0}{V_0^2} } \Big)"


We take the positive root and we calculate the total time with V0 = 15 m/s and h0 = 50 m:


"t= (\frac{15\,m/s}{9.8\,m/s^2})\Big( \sqrt{ 1+\frac{2(9.8\,m/s^2)(50\,m)}{(15\,m/s)^2} } +1\Big)
\\ t= ( 1.53\,s)(1+\sqrt{ 1+4.355 })= 5.07\,s"


b) the total distance along the y axis will be twice the maximum distance reached when the ball goes upwards and then we add it to the height of the cliff h0 to find the total distance traveled by the ball as D:


"t_{upwards}= \frac{V_0}{g}
\\ d=V_0t-\frac{gt^2}{2}=V_0(\frac{V_0}{g})-\frac{g}{2}(\frac{V_0}{g})^2
\\ \therefore d = \frac{V_0^2}{g}-\frac{V_0^2}{2g}=\frac{V_0^2}{2g}
\\ \to D = h_0 + 2d = h_0 +\frac{V_0^2}{g} 
\\ D= 50\,m+\frac{[15\frac{m}{s}]^2}{9.8\frac{m}{s^2}}= 72.96\,m"


In conclusion, a) it takes 5.07 s for the ball to reach the base of the cliff, and b) the ball covers a total distance of 72.96 m.